FRS Home Page
-
–
S2
–
Pictures
–
Drawings
–
Cable Connections –
TPC Electronics – Memo for preparation before beam time
Official Manual of TPC (
PDF )
Recent Log and Memo about TPC (2019) (External link)
Present Configuration (as of 2020/Oct/22)
- in a box (pocket) for the S2 vacuum chamber (GTS3EG7T)
- Window = 100 um Stainless steel (each of front and behind)
- Gap = 8 cm (drift HV = -3200 V)
- Normal orientation
- Now calibration positions should be x = -12, 0, +12 mm and y = -6, 0, +6 mm
- To have consistent track from all 4 TPCs, one could add the values on page 5 of this memo to the grid position
- Maximum allowed HV: -3200 V (drift), +1100 V (anode), -700 V (calibration, PMT).
- For heavy ions, anode and PMT voltages should be much lower than the maximum values.
Status during previous beam time
2021/Feb/15-Feb/22 (12C beam, S533/BARB + EXPERT test)
- HV(anode) = +1050V, HV(drift) = -3200V, HV(grid-pmt) = -700V
- Analog raw signal around 150--200 mV. Vth(tpc module) around 45--60 mV
- Calibration file is /d/land4/sec/s533/tpc22_calib_0024.lmd (Logbook p41-43)
- Parameter sets used during experiment is found
here.
2020/June/05--June/08 (proton beam, WASA000)
- Applied voltages: anode HV = +1150 V, drift = -3200 V, calib-pmt =
- Calibration file: wasa2020june_tpc22calib_[0015-0018].lmd
2020/May/xx--May/xx (Pb beam, S469, very low evenry)
2020/Apr/08--Apr/27 (Pb beam, S468, S482, S481)
- Applied voltages: anode HV = +580 V, drift HV = -3200V, fiber-PMT = -350V
2020/Feb/28--Mar/16 (Xe beam, S474, S459, S480)
- Applied voltages: anode HV = +600 V, drift HV = -3200V, fiber-PMT = -350V
- Used very short time to get the first impression about the beam poisition. Calibration grid (run0114).
Parameters - There are files with all TPC21+22+23+24, which can be used for overall check.
2020/Feb/20--Feb/26 (Kr beam, Ca fragments, Cave-C S467)
- Applied voltages: anode HV = +700 V, drift HV = -3200V.
- Used very short time to get the first impression about the beam poisition. No grid caliberation was performed.
- There are files with all TPC21+22+23+24, which can be used for calibration (if needed). Runs 92--95.
2019/Dec/09--Dec/19 (U beam)
- After changing primary beam from 40Ar to 238U.
- Applied voltages: anode HV = +550 V, drift = -3200V
2019/Nov/18--Dec/06 (Ar beam)
- Installed at S2 as TPC22.
- All signals were checked and ok, calibration scintllator double peak ?? >> to be investigated.
- Gap is 8cm. normal orientation. DL1 signal is fine (see comment in 2019-April run).
- Applied voltages: anode HV = +800 V, drift = -3200V, fiber PMT = -500V.
2019/April/1--3 (107Ag, 290MeV/u, primary beam)
- Installed at S2 as TPC21.
- Signal of TPC22-DL1 is abnormally small. Swapping cables did not work. Something in the TPC chamber should be the reason.
- Other signals are reasonable.
- Calibration scintllator was ok.
- normal orientation - We need to open this TPC after the beam time to investigate problems of DL1.
--
NisanYaprakBenal - 2026-10-26
Updated November 07, 2019
Messagelog
--
SagarRoy - 2021-06-26